OpenStudy (anonymous):

Complete the general form of the equation of a sinusoidal function having an amplitude of 1, a period of pi/2, and a vertical shift up 3 units.

OpenStudy (anonymous):

general form

OpenStudy (anonymous):

thank you

OpenStudy (perl):

The general sin x is y = A sin [2*Pi/per.*( x- p.s.)] + v.s. A = amplitude per. = period p.s. = phase shift or horizontal shift v.s. = vertical shift

OpenStudy (anonymous):

Do you know the answer?

OpenStudy (perl):

now you can substitute :)

OpenStudy (perl):

amplitude of 1, a period of pi/2, and a vertical shift up 3 units. y = A sin [2*Pi/per.*( x- p.s.)] + v.s. y = 1 sin [2pi / (pi/2) *( x - 0) ] + 3

OpenStudy (perl):

and you can simplify that

OpenStudy (anonymous):

y= 1sin [2 pi/ pi/2] (x-0)] +3

OpenStudy (anonymous):

yeah thank u

OpenStudy (perl):

right, and 2pi / (pi/2) = 2pi * 2/pi = 4